# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ III.1.4 Monolayer 2H- and 1T-VSe2
Monolayer VSe2 has been a source of controversy throughout the theoretical and experimental literature, with substantial claims of near-room temperature ferromagnetism (from 291 K to 470 K [149, 191, 192, 193]).The T (octahedral phase (1T)-centered honeycombs) phase and the H (the trigonal prismatic phase (2H)-hexagonal honeycombs) phase (shown in Fig. 4) of 2D VSe2 have a close lattice match and similar total energies, which makes it difficult to discern which phase is being observed experimentally [194, 195, 191, 149]. The difficulty discerning the relative stability of the phases and related structural uncertainty undoubtedly imply that the structural parameters are coupled to the magnetic properties [149, 191, 192, 193, 196]. Similarly to Ref. 109 and Ref. 175, we used a combination of DMC, DFT, and the surrogate Hessian line-search structural optimization technique [174] to clarify the discrepancies in structural parameters and relative phase stability of the T and H phases of monolayer VSe2.
Fig. 4 depicts a summary of DFT results (multiple functionals with and without U) alongside DMC (red bars) for the lattice constant (), V-Se bond distance (), and the relative phase energy (E). As expected, there is a large disagreement between DFT functionals for , , and E, indicating the need for a theory such as DMC. With DMC and the Hessian line-search method, we computed the lattice constant and V-Se distance to be 3.414(12) Å and 2.505(7) Å , respectively, for T-VSe2 and 3.335(8) Å and 2.503(5) Å , respectively, for H-VSe2. The relative energy between the T and H phases was found to be 0.06(2) eV, with the H phase being lower in energy than the T phase in freestanding form. We then constructed a phase diagram between T and H with DMC accuracy and found that a H-to-T phase transition can be induced by applying small amounts of strain ( 3 ). More details of this study can be found in Ref. 176. As a follow-up to this work, the authors intend to focus on studying the magnetic properties of 2D T-VSe2 with DMC methods, specifically running DMC for various magnetic configurations and predicting the transition temperature.
@@ III.2 Electronic Properties
Electronic structure: It is now established that 2D systems such as graphene, phosphorene, and transition metal dichalcogenides belong to a class of materials that is exceptionally interesting for both fundamental research and for future use in technology. The electronic structure of 2D systems is astonishingly rich, since it provides an unusual combination of periodicity and free boundaries, the possibility of stacking with new variational freedom (twistronics), significant effects related to the presence of substrates, and further potential modifications with doping, straining (straintronics), and related processing. Due to the presence of unique electronic phases, these systems pose a number of challenges for any theoretical or computational method.
@@
It is useful to specify that, from the point of view of electronic structure theory, the 2D materials in this review are not strictly systems with electrons in a 2D plane, but rather 2D slabs with atomic thickness. Consequently, the electronic structure, one-particle orbitals, and other properties indeed depend on all three spatial coordinates. (Previous theoretical work has shown that the calculations of slabs have their intricacies, as can be seen for the homogeneous electron gas in a slab geometry, see Ref. 197.) This is important since many properties that apply to strictly 2D geometries might not apply or might be significantly modified. Clearly, electrons have a nonzero probability to be on either side of the material in the orthogonal direction, and they are not confined within a 2D plane. This leads to the possibility of Janus materials ,[198] in which the opposing surfaces have different properties due to different functionalizations, doping, and symmetry-breaking. An important consequence of this may be seen in the electronic structure, as most clearly revealed by the dependence of the dielectric constant on the inter-electronic distance. In 3D solids, the dielectric constant grows monotonically until it saturates at large distances to the bulk value. In 2D materials, it initially grows, however, at the distance of a few atomic bonds, it exhibits a rather sharp maximum and then rapidly falls off to a small value. This small value corresponds to long-range distance vdW-like behavior where distant regions appear electrically neutral with typical dipole or higher order dispersive fluctuations. This implies that, at short distances, we see the effects of bonds which look like locally 3D material while at large distances the system exhibits dispersive effects resembling effective interactions with non-covalent character.
@@
DFT and GW methods have been widely used to study the electronic structure of 2D systems. The accuracy of the former is determined by the XC-functional, whereas that of the latter by the unperturbed state and the way the perturbation expansion is terminated. This introduces significant biases between the two approaches as well as within each method. Quasiparticle band gaps typically between these methods by up to 1 eV. As a result, experiments are expected to provide the final answer. However, 2D materials typically need to be supported by substrates (quartz, sapphire, etc.) which compromises their 2D character by providing unintentional dielectric embedding. As shown in Fig. 5, the experimental effect of the tuning by the substrate is also of the order of 1 eV, complicating the comparison with calculations. In such a situation, QMC may provide the final answers and even challenge the achievable experimental accuracy. The situation is illustrated in Fig. 6 where the quasiparticle gap of monolayer phosphorene calculated by DFT (PBE and B3LYP), GW (G0W0 and GW0 using both PBE and B3LYP wavefunctions), and QMC methods is compared to the rare result where the experimental band gap is measured on a freestanding sample [199]. The QMC result is seen to agree with the experimental result almost with chemical accuracy, given that no vibronic effects were considered in the QMC calculation, while both the DFT and GW methods tend to underestimate the band gap and can feature a very wide spread in their calculated values.
@@
Excited states: One of the key characteristics that is important for applications is the value of the band gap and its sensitivity to variations that come with the impact of doping, substrates, strain, multilayer stacking, and both internal (e.g., spin-orbit) and external influences. In QMC calculations of band gaps the key complications come from two main sources of error, namely, finite-size and fixed-node biases. The fixed-node error is determined by the quality of the antisymmetric part of the trial function which is often just single-reference since robust methodologies for multi-reference wavefunctions for periodic systems are under development. In QMC the band gaps can be calculated essentially by two approaches as outlined in the following.
@@
-Fundamental band gap: One way is to use the definition of the fundamental gap which implies charging the system with electrons by an additional electron , then by an additional hole and calculating the response as the difference of total energies for charged and neutral states. This has been routinely applied in 3D and often also to a number of 2D systems [203, 204, 175, 205]. Some cases showed a minor (0.1 to 0.2) eV upward bias that can be qualitatively understood by the fact that cation typically relaxes more than the anion since the anionic state with its conduction state occupation should typically involve more than a single-configuration. Hence, the resulting minor bias which is, in fact, surprisingly small considering just a single-reference approximation of the nodal hypersurfaces for both charged and neutral systems.
@@
We note that additional bias could come from finite size scaling since the charged states have to be compensated in order to make the periodic supercell neutral. The commonly used compensation by the opposite sign constant background charge eliminates the leading (monopole) diverging term, however, it does not perfectly compensate the subleading terms and although the impact mostly appears to be small [206], this need not be valid. There are also subtleties in using charge compensation and Ewald sums when treating 2D systems and increasing size of box in the orthogonal direction. An example for band gap in hexagonal Boron Nitride (hBN) [207] is shown in Fig. 7 where first the in-plane finite-size scaling is shown for various lengths of the simulation cell vector orthogonal to 2D material (Lz) with the inset showing the Lz convergence of the in-plane scaled band gaps. It is seen that as the neutralizing background dilutes with increasing Lz, the band gap appears to diverge with Lz due to artificial charge decompensation. In fact, proper compensation for slab geometries in real systems is still a matter of current studies [207]. In addition to our own work, Hunt et al. [205] previously studied the electronic and excitonic properties of 2D hBN with DMC, carefully taking finite-size errors and vibrational renormalization into account.
@@
-Promotion band gap: The second method is closer to experiments that provide optical spectrum, namely, promotion of an electron from valence into the conduction band. This excitation can be direct or indirect (i.e., approximately modelling a phonon-assisted optical excitation). Since the optical spectrum typically involves excitonic states, it is good to distinguish the two possibilities that can occur:
@@
Wannier excitons: In systems with strong covalent bonds and large dielectric constants, one typically encounters Wannier excitons with exciton binding energies from meVs to a few tens of meV and therefore the effect on the electronic bands is very small. This is what is observed in many such calculations[206] since any relaxation into the excitonic state provides only very minor effects which are at the level of statistical error bars in DMC.
@@
Frenkel excitons: For systems with possible presence of excitons with binding energies larger than 0.5 eV (Frenkel excitons) the situation is more complicated.
@@
(i) In covalent or mainly covalent materials, the one-particle orbitals from band structure calculations, such as DFT, HF, DFT+U, etc., provide Bloch orbitals that enter the trial wavefunction. This determines the symmetry and periodicity of the resulting trial wavefunction. In most cases, this constraint "locks" the overall band structure picture also in QMC calculations. Typical QMC effects come as band gap shifts, however, the overall structure of the bands remains pretty much the same. Therefore, due to the nodal constraint, the relaxation to more localized, more strongly bonded non-Bloch excitonic state is strongly hindered. This is easy to understand since the localized exciton wavefunction in Bloch state basis necessarily requires large multi-reference expansion. Therefore, any relaxation to an excitonic state is mostly absent [201, 203], which shows that the overlap of the exciton and Bloch state promotion is negligible or small. Note that in the non-interacting limit the overlap identically vanishes showing thus the dominant one-particle nature of this behavior.
@@
(ii) However, in some cases the nodal constraint might be less restricting. One such possibility could occur if the excitonic state and Bloch excitation share a subgroup of symmetries. Possible mixing of such two states would have to be carried out explicitly. Another possibility could occur in environments with rapidly changing charge densities (such as ionic bonds or 2D layered materials with large space between the layers) since absence of proper tails in the trial function can hinder the needed charge redistribution. Therefore, the charge relaxation and/or localization could be compromised with resulting energy decrease towards the excitonic state. This could be observed in some systems [175], although the signal is comparable to the error bar range due to complications with thermodynamic limit extrapolation. It is clear that this requires a new leap in quality of the trial function in order to quantify the involved effects. There is no fundamental obstacle in carrying out such analysis, the issues are mainly technical due to limited availability of appropriate software tools. Clearly, one would need to rebuild the trial function with localized orbitals, possibly orthogonalize it to the Bloch excited state that corresponds to the fundamental gap excitation, to make the distinction of fundamental vs. excitonic state explicitly. Note that is this routinely done in molecular studies since many quantum chemical codes enable to carry out corresponding calculations. Another such case could occur for vdW molecular crystals, where the localized state(s) on a single molecule could dominate the low-lying excitation spectrum since the periodicity causes only minor energetic shifts and results in very flat low-lying bands. More dispersive bands appear in significantly higher scattering states in the conduction part of the spectrum.
@@
Several QMC works have computed the fundamental and promotion band gap to estimate the exciton binding energy of 2D systems [203, 204, 175, 205]. Additionally, there have been studies by Szyniszewski et al. [208] and Mostaani et al. [209] that have utilized DMC coupled with other theoretical approaches to successfully estimate biexciton and trion energies in 2D semiconductors.
@@ III.2.1 Monolayer phosphorene and MoS2
Alongside graphene, phosphorene and transition metal dichalcogenides belong to the most studied 2D semiconductors. Their band gap is often direct. In phosphorene the band gap is direct at in a monolayer and few-layers and direct even in bulk black phosphorus albeit at the Z point. In MoS2 it is direct at the K-point only in the monolayer. In addition, 2D semiconductors usually possess ultrahigh carrier mobility and field-effect switching ratios, which make them ideal materials for field-effect and digital logic transistors. 2D semiconductors with appropriate and tuneable direct band gap have achieved extremely efficient photon absorption, emission, and photoelectric conversion and have been widely used in the field of optoelectronic devices. To meet more demands, various techniques have been exploited to modulate their properties, including doping, alloying, forming vdW heterostructures, and strain engineering orstraintronics [210, 211]. Due to their atomic thickness, 2D materials are highly sensitive to external perturbations, such as strain. Their resilience to mechanical deformations allows application of strains well in excess of 10 %. By applying strain, the lattice and electronic structure is modulated as well as their various properties, such as the carrier mobility. MoS2 has traditionally been considered the quintessential straintronic material for which many straintronic experiments have been performed [212].
@@
Recently, the straintronic response of two 2D semiconductors, monolayer phosphorene and MoS2 has been studied using fixed-node QMC methods [202, 213]. In phosphorene the strain was applied in both armchair and zigzag direction by adjusting the a and b lattice parameters, see Fig. 8, considering deformations of up to 10 %. Determination of strained properties was treated as a full optimization problem in the space of four structural variables: lattice parameters a, b and two internal parameters x, y[202]. Around the minima the ground-state potential energy surface (PES) was fitted by 4D paraboloid functions used to find the lowest point on the x, y subspace, leaving to further minimize only bivariate functions . The excited state was computed only at the minimum for the internal parameters x, y and the quasiparticle band gap for any applied strain was computed as , subject to finite-size scaling. In MoS2 only diagonal strain was considered and a scaled down version to two parameters (one lattice parameter a and one internal parameter x) was used leaving a single function to optimize [213]. This approach obviously allows to study the 2D materials both in- and out-of equilibrium.
@@
For phosphorene in equilibrium, we obtain (a = 3.30 0.003 Å), (b = 4.61 0.006 Å), and ( = 2.53 0.020 eV), see also Fig. 6. This value of the quasiparticle gap is in excellent agreement with the experimental value for freestanding monolayer phosphorene of 2.46 eV [199], see also Fig. 6 (keeping in mind that neglecting adiabatic, vibronic, and zero-point vibrational energy tends to increase the gap value compared to experiments). The QMC optimized structure exhibits noticeable differences with respect to the structure derived from the 3D black phosphorus crystal especially in the b parameter (4.376Å). This variation of the b parameter in 3D crystal is due to partially chemical interlayer interaction which reduces it by 5 % and is almost completely absent from MoS2, where the interlayer interaction in the 3D crystal is fully vdW [213]. These trends are fairly well described by the DFT-PBE functional and not as well by the hybrid functionals [202]. All gaps in the DFT treatment are appreciably smaller than the QMC value. As expected, the smallest value by about 2 eV is obtained by the DFT-PBE functional. The hybrid functionals yield larger values but fail in predicting the equilibrium geometries.
@@
The calculated band gap phase diagram for phosphorene is shown in Fig. 8. The diagram was constructed by calculating the boundaries between the and a band gap at formed by interchange of LUMO (Least unoccupied molecular orbital) and LUMO + 1 states () and the boundary between the and X due to applied strain from the intersections of the respective PESs. As for the first boundary, while both gaps are direct at , the nature of the excited state is different if the order of unoccupied states is interchanged. The new LUMO state has a differing curvature, hence, the transport properties in the conduction band are expected to be significantly modified which goes in line with straintronic as a tool for electrom effective mass modification. The QMC boundaries outline a strain tuning area for the direct band gap more than twice as large as that determined by DFT-PBE (and similarly by hybrid functionals [202]). This area extends mostly into the region of tensile strain which is important to prevent the material from wrinkling at higher compressive loads. The DFT results are qualitatively similar, see Fig. 8, irrespective of the DFT functional. All PESs (QMC, DFT) are rather parallel, alebeit strongly shifted vertically in energy and their intersections all feature similar curvatures. The primary differences consist of relative band offsets (magnitude of band gaps) and the location of the minima. These two factors play havoc with the form of the phase diagrams. In phosphorene, Fig. 8, the offsets reduced the section of the diagram by 50 % in phosphorene [202]. In MoS2, where the strain tuning areas are more compressed, the different treatments were predicting qualitatively different tunings in the KK section: QMC mainly in the compressiveregion, GW and B3LYP exclusively in compressive, whereas HSE mainly into tensile [213]. Phosphorene was also found to feature high values of the gauge factor of the order of 100 meV per %, which is comparable to MoS2. Combined with the huge tuning area of several percent leaving the direct nature of the band gap intact, this provides a colossal straintronic tunability of phosphorene with the tuning limit likely determined only by the mechanical breakdown of the material. The range of band-gap tuning by applied strain while maintaining the direct band gap at is truly huge with achievable values of the band gap in the range 2.1 eV to 3.8 eV. Surprisingly, this is at complete variance with the straintronic response of the quintessential straintronic material, monolayer MoS2, which features similar gauge factor but an order of magnitude smaller tuning area [213].

### v2

@@ III.1.4 Monolayer 2H- and 1T-VSe2
Monolayer VSe2 has been a source of controversy throughout the theoretical and experimental literature, with substantial claims of near–room temperature ferromagnetism (from 291 K to 470 K [184, 223, 224, 225]).The T (octahedral phase (1T)-centered honeycombs) phase and the H (the trigonal prismatic phase (2H)-hexagonal honeycombs) phase (shown in Fig. 3) of 2D VSe2 have a close lattice match and similar total energies, which makes it difficult to discern which phase is being observed experimentally.[226, 227, 223, 184] The difficulty in discerning the relative stability of the phases and related structural uncertainty undoubtedly imply that the structural parameters are coupled to the magnetic properties.[184, 223, 224, 225, 228] Similarly to Refs. 109 and 146, we used a combination of DMC, DFT, and the surrogate Hessian line-search structural optimization technique [209] to clarify the discrepancies in structural parameters and relative phase stability of the T and H phases of monolayer VSe2.
Fig. 3 depicts a summary of DFT results (multiple functionals with and without U) alongside DMC (red bars) for the lattice constant (), V-Se bond distance (), and the relative phase energy (E). As expected, there is a large disagreement among DFT functionals for , , and E, indicating the need for a theory such as DMC. With DMC and the Hessian line-search method, we computed the lattice constant and V-Se distance to be 3.414(12) Å and 2.505(7) Å, respectively, for T-VSe2 and 3.335(8) Å and 2.503(5) Å, respectively, for H-VSe2. The relative energy between the T and H phases was found to be 0.06(2) eV, with the H phase being lower in energy than the T phase in freestanding form. We then constructed a phase diagram between the T and H phases with DMC accuracy and found that an H-to-T phase transition can be induced by applying small amounts of strain (3). More details of this study can be found in Ref. 174. As a follow-up to this work, the authors intend to focus on studying the magnetic properties of 2D T-VSe2 with DMC methods, specifically running DMC for various magnetic configurations and predicting the transition temperature.
@@ III.2 Electronic Properties
Electronic structure: It is now established that 2D systems such as graphene, phosphorene, and TMDs belong to a class of materials that is exceptionally interesting for both fundamental research and future technological use. The electronic structure of 2D systems is astonishingly rich because it provides an unusual combination of periodicity and free boundaries, the possibility of stacking with new variational freedom (twistronics), significant effects related to the presence of substrates, and further potential modifications with doping, straining (straintronics), and related processing. Due to the presence of unique electronic phases, these systems pose a number of challenges for any theoretical or computational method.
@@
From the perspective of electronic structure theory, the 2D materials in this review are not strictly systems with electrons in a 2D plane but rather 2D slabs with atomic thickness. Consequently, the electronic structure, one-particle orbitals, and other properties indeed depend on all three spatial coordinates. (Previous theoretical work has shown that calculations of slabs have their intricacies, as can be seen for the homogeneous electron gas in a slab geometry; see Ref. 229.) This is important because many properties that apply to strictly 2D geometries might not apply or might be significantly modified. Clearly, electrons have a nonzero probability to be on either side of the material in the orthogonal -direction, and they are not confined within a 2D plane. This leads to the possibility of Janus materials,[230] in which the opposing surfaces have different properties due to different functionalizations, doping, and symmetry-breaking. An important consequence of this may be seen in their electronic structure, as most clearly revealed by the dependence of their dielectric constant on the interelectronic distance. In 3D solids, the dielectric constant grows monotonically until it saturates at large distances to the bulk value. In 2D materials, it initially grows; however, at the distance of a few atomic bonds, it exhibits a rather sharp maximum and then rapidly falls off to a small value. This small value corresponds to long-range-distance vdW-like behavior in which distant regions appear electrically neutral with typical dipole or higher-order dispersive fluctuations. This implies that at short distances, the system’s bonds look locally like those in a 3D material, while at large distances, the system exhibits dispersive effects resembling effective interactions with noncovalent character.
@@
The DFT and GW methods have been widely used to study the electronic structure of 2D systems. The accuracy of the former is determined by the XC functional, whereas that of the latter is determined by the unperturbed state and the way the perturbation expansion is terminated. This introduces significant biases between the two approaches as well as within each method. Quasiparticle band gaps typically differ between these methods by up to 1 eV. As a result, experiments are expected to provide the final answer. However, 2D materials typically must be supported by substrates (quartz, sapphire, etc.), which compromises their 2D character by providing unintentional dielectric embedding. As shown in Fig. 4, the experimental effect of the tuning by the substrate is also of the order of 1 eV, complicating the comparison with calculations. In such a situation, DMC may provide the final answers and even challenge the achievable experimental accuracy. The situation is illustrated in Fig. 5, where the quasiparticle gap of monolayer phosphorene calculated by the DFT (PBE and B3LYP), GW (G0W0 and GW0 using both PBE and B3LYP wavefunctions), and DMC methods is compared to the rare result where the experimental band gap is measured on a freestanding sample.[231] The DMC result agrees with the experimental result almost with chemical accuracy, given that no vibronic effects were considered in the DMC calculation, while both the DFT and GW methods tend to underestimate the band gap and can feature a very wide spread in their calculated values.
@@
Excited states: One of the key characteristics of a 2D material important for many applications is the value of the band gap and its sensitivity to variations caused by doping, substrates, strain, multilayer stacking, and both internal (e.g., spin-orbit) and external influences. In DMC calculations of band gaps, the key complications come from two main sources of error—namely, FS and fixed-node biases. The fixed-node error is determined by the quality of the antisymmetric part of the trial function, which is often just single reference because robust methodologies for multireference wavefunctions for periodic systems are under development. In QMC, the band gaps can be calculated essentially by two approaches as outlined in the following.
@@
Fundamental band gap: One way is to use the definition of the fundamental gap, which implies charging the system with electrons by an additional electron and then by an additional hole and calculating the response as the difference of total energies for charged and neutral states. This has been routinely applied in 3D and often also to a number of 2D systems.[173, 232, 146, 148] Some cases showed a minor (0.1 to 0.2) eV upward bias that can be qualitatively understood by the fact that the cation typically relaxes more than the anion because the anionic state with its conduction state occupation should typically involve more than a single configuration. The reason is that the conduction state excitation could mix other available (i.e., unoccupied) conduction states that are not too far energy wise and that can form linear combinations that produce the same symmetry as in the excitation. It is possible to have either increases or decreases in energy from mixing of excitations and each case requires careful analysis. It is even more pertinent when energy differences (band gaps) are involved, since each of the corresponding states can be impacted differently. In practice, the conduction states are often less accurate and less well converged than the ground state calculation. This explains the resulting minor bias, which is surprisingly small considering just a single-reference approximation of the nodal hypersurfaces for both charged and neutral systems.
@@
Additional bias could come from FS scaling because the charged states must be compensated to make the periodic supercell neutral. The commonly used compensation by the opposite sign constant background charge eliminates the leading (monopole) diverging term. However, it does not perfectly compensate the subleading terms, and although the impact mostly appears to be small,[144] this need not be valid. There are also subtleties in using charge compensation and Ewald sums when treating 2D systems and increasing the size of the box in the orthogonal direction. An example for band gap in hBN[233] is shown in Fig. 6, where first the in-plane FS scaling with the number of primitive unit cells, , is shown for various lengths of the simulation cell vector orthogonal to 2D material (Lz). The inset shows the Lz convergence of the in-plane scaled band gaps. As the neutralizing background dilutes with increasing Lz, the band gap appears to diverge with Lz due to artificial charge decompensation. In fact, proper compensation for slab geometries in real systems is still a matter of current studies.[233] In addition to our own benchmarking calculations displayed in Fig. 6, Hunt et al. [148] previously studied the electronic and excitonic properties of 2D hBN with DMC, carefully taking FS errors and vibrational renormalization into account (discussed in more detail in Section III.2.3).
@@
Promotion band gap: The second method is closer to experiments that provide the optical spectrum—namely, promotion of an electron from valence into the conduction band. This excitation can be direct or indirect (i.e., approximately modeling a phonon-assisted optical excitation). Because the optical spectrum typically involves excitonic states, it is good to distinguish the two possibilities that can occur:
@@
Wannier excitons: In systems with strong covalent bonds and large dielectric constants, one typically encounters Wannier excitons with exciton binding energies from meVs to a few tens of meV, so the effect on the electronic bands is very small. This is what is observed in many such calculations[144] because any relaxation into the excitonic state provides only very minor effects, which are at the level of statistical error bars in DMC.
@@
Frenkel excitons: For systems with a possible presence of excitons with binding energies larger than 0.5 eV (Frenkel excitons), the situation is more complicated.
@@
(i) In covalent or mainly covalent materials, the one-particle orbitals from band structure calculations, such as DFT, HF, and DFT+U, provide Bloch orbitals that enter the trial wavefunction. This determines the symmetry and periodicity of the resulting trial wavefunction. In most cases, this constraint “locks” the overall band structure picture also in QMC calculations. Typical QMC effects come as band gap shifts; however, the overall structure of the bands remains nearly the same. Therefore, due to the nodal constraint, the relaxation to a more-localized, more strongly bonded non-Bloch excitonic state is strongly hindered. This is easy to understand because the localized exciton wavefunction in the Bloch state basis necessarily requires large multireference expansion. Therefore, any relaxation to an excitonic state is mostly absent, [149, 173] which shows that the overlap of the exciton and Bloch state promotion is negligible or small. In the noninteracting limit, the overlap identically vanishes, thus showing the dominant one-particle nature of this behavior.
@@
(ii) However, in some cases, the nodal constraint might be less restricting. One such possibility could occur if the excitonic state and Bloch excitation share a subgroup of symmetries. Possible mixing of such two states would have to be carried out explicitly. Another possibility could occur in environments with rapidly changing charge densities (e.g., ionic bonds or 2D layered materials with large space between the layers) because absence of proper tails in the trial function can hinder the needed charge redistribution. Therefore, the charge relaxation and/or localization could be compromised with a resulting energy decrease toward the excitonic state. This could be observed in some systems, [146] although the signal is comparable to the error bar range due to complications with thermodynamic limit extrapolation. This requires a new leap in the quality of the trial function to quantify the involved effects. There is no fundamental obstacle in carrying out such analysis; the issues are mainly technical due to limited availability of appropriate software tools. Clearly, one would need to rebuild the trial function with localized orbitals and possibly orthogonalize it to the Bloch excited state that corresponds to the fundamental gap excitation to distinguish the fundamental vs. excitonic state explicitly. This is routinely done in molecular studies because many quantum chemical codes enable corresponding calculations. Another such case could occur for vdW molecular crystals, where the localized state(s) on a single molecule could dominate the low-lying excitation spectrum because the periodicity causes only minor energetic shifts and results in very flat low-lying bands. More dispersive bands appear in significantly higher scattering states in the conduction part of the spectrum.
@@
Several QMC works have computed the fundamental and promotion band gap to estimate the exciton binding energy of 2D systems.[173, 232, 146, 148] Additionally, there have been studies by Szyniszewski et al. [234] and Mostaani et al. [235] that have utilized DMC coupled with other theoretical approaches to successfully estimate biexciton and trion energies in 2D semiconductors.
@@ III.2.1 Monolayer Phosphorene and MoS2
Alongside graphene, phosphorene and TMDs belong to the most studied 2D semiconductors. Their band gap is often direct. In phosphorene, the band gap is direct at in a monolayer and few-layers and direct even in bulk black phosphorus, albeit at the Z point. In MoS2, it is direct at the K point only in the monolayer. In addition, 2D semiconductors usually possess ultrahigh carrier mobility and field-effect switching ratios, which make them ideal materials for field-effect and digital logic transistors. 2D semiconductors with an appropriate and tunable direct band gap have achieved extremely efficient photon absorption, emission, and photoelectric conversion and have been widely used in the field of optoelectronic devices. To meet more demands, various techniques have been exploited to modulate their properties, including doping, alloying, forming vdW heterostructures, and strain engineering or straintronics.[236, 237] Due to their atomic thickness, 2D materials are highly sensitive to external perturbations, such as strain. Their resilience to mechanical deformations allows application of strains well in excess of 10%. Applying strain enables the lattice and electronic structure to be modulated as well as their various properties, such as the carrier mobility. MoS2 has traditionally been considered the quintessential straintronic material for which many straintronic experiments have been performed.[238]
@@
Recently, the straintronic response of two 2D semiconductors, monolayer phosphorene and MoS2, has been studied using fixed-node DMC methods.[172, 239] In phosphorene, the strain was applied in both armchair and zigzag directions by adjusting the a and b lattice parameters (Fig. 7), considering deformations of up to 10%. These deformations determine the in-plane , strains. Determination of strained properties was treated as a full optimization problem in the space of four structural variables: lattice parameters a, b and two internal parameters x, y.[172] Around the minima, the ground-state PES was fitted by 4D paraboloid functions used to find the lowest point on the x, y subspace, leaving to further minimize only bivariate functions . The excited state was computed only at the minimum for the internal parameters x, y, and the quasiparticle band gap for any applied strain was computed as , subject to FS scaling. In MoS2, only diagonal strain was considered, and a scaled-down version to two parameters (one lattice parameter a and one internal parameter x) was used, leaving a single variable function to optimize.[239] This approach allows researchers to study the 2D materials both in and out of equilibrium.
@@
For phosphorene in equilibrium, we obtain (a = 3.30 0.003 Å), (b = 4.61 0.006 Å), and ( = 2.53 0.020 eV); see also Fig. 5. This value of the quasiparticle gap is in excellent agreement with the experimental value for freestanding monolayer phosphorene of 2.46 eV;[231] see also Fig. 5 (keeping in mind that neglecting adiabatic, vibronic, and zero-point vibrational energy tends to increase the gap value compared to experiments). The DMC optimized structure exhibits noticeable differences with respect to the structure derived from the 3D black phosphorus crystal especially in the b parameter (4.376Å). This variation of the b parameter in 3D crystal is partially due to chemical interlayer interaction that reduces it by 5% and is almost completely absent from MoS2, where the interlayer interaction in the 3D crystal is fully vdW.[239] These trends are fairly well described by the DFT-PBE functional and not as well by the hybrid functionals.[172] All gaps in the DFT treatment are appreciably smaller than the DMC value. As expected, the smallest value by about 2 eV is obtained by the DFT-PBE functional. The hybrid functionals yield larger values but fail in predicting the equilibrium geometries.
@@
The calculated band gap phase diagram for phosphorene is shown in Fig. 7. The diagram was constructed from intersections of PESs corresponding to conditions under which the system retains the zero-strain direct gap, and where it forms a direct gap at the point but with a level crossing between the lowest unoccupied molecular orbital (LUMO) and LUMO+1 states, labeled , and a third situation with a X indirect gap. The three different PESs outline three different “phase” boundaries. No attempt was made to determine the other phase boundaries in Fig. 7; hence, they correspond to DFT-PBE boundaries. As for the first boundary, while both gaps are direct at , the nature of the excited state is different if the order of unoccupied states is interchanged. The new LUMO state has a differing curvature; hence, the transport properties in the conduction band are expected to be significantly modified, which goes in line with straintronic as a tool for electron effective mass modification. The DMC boundaries outline a strain tuning area for the direct band gap more than twice as large as that determined by DFT-PBE (and similarly by hybrid functionals[172]). This area extends mostly into the region of tensile strain, which is important to prevent the material from wrinkling at higher compressive loads. The DFT results are qualitatively similar (Fig. 7) regardless of the DFT functional. All PESs (e.g., QMC, DFT) are rather parallel, albeit strongly shifted vertically in energy, and their intersections all feature similar curvatures. The primary differences consist of relative band offsets (magnitude of band gaps) and the location of the minima. These two factors play havoc with the form of the phase diagrams. In phosphorene (Fig. 7), the offsets reduced the section of the diagram by 50% in phosphorene.[172] In MoS2, where the strain tuning areas are more compressed, the different treatments were predicting qualitatively different tunings in the KK section: DMC mainly in the compressive region, GW and B3LYP exclusively in compressive, and HSE mainly into tensile.[239] Phosphorene was also found to feature high values of the gauge factor of the order of 100 meV per %, which is comparable to MoS2. Combined with the huge tuning area of several percent leaving the direct nature of the band gap intact, this provides a colossal straintronic tunability of phosphorene with the tuning limit likely determined only by the mechanical breakdown of the material. The range of band-gap tuning by applied strain while maintaining the direct band gap at is truly huge with achievable values of the band gap in the range 2.1 eV to 3.8 eV. Surprisingly, this is at complete variance with the straintronic response of the quintessential straintronic material, monolayer MoS2, which features similar a gauge factor but an order-of-magnitude smaller tuning area.[239]